Evaluate the expression for the given value of x and y:
2x^2-1
______; x= - 1/2, y=2/3 Im lost and don't know were to start
y+3

Answer by funmath(2933)

(2x^2-1)
________;x=-1/2, y=2/3
(y+3)
=(2(-1/2)^2-1)
____________
(2/3+3)
=(2(-1/2)(-1/2)-1)
__________________
(2/3+3)
=(2(1/4)-1)
_____________
(2/3+3)
=(1/2-1)
__________;LCD=6, so multiply top and bottom by 6.
(2/3+3)
=6(1/2-1)
_________
6(2/3+3)
=(6/2-6)
_________
(12/3+18)
=(3-6)
_________
(4+18)
=-3/22